OpenStudy (anonymous):

Check if I'm right? 2(x-y) < -5 And from that I got.... Y < 1x + 5/2

jimthompson5910 (jim_thompson5910):

Did you mean to write \[\Large y < 1x + \frac{5}{2}\] ?

OpenStudy (anonymous):

Yes.

jimthompson5910 (jim_thompson5910):

If so, then you are close, very close, but remember that the sign will flip when you multiply or divide both sides by a negative number

jimthompson5910 (jim_thompson5910):

ex: -2x > 4 turns into x < -2

OpenStudy (anonymous):

That's incorrect. When you multiply both sides by -1, the face of the inequality switches. \[2x-2y < -5\]

OpenStudy (anonymous):

Even if the sign is flipped, I don't see how I can be right because it tells me to graph and the graph only goes up to 5.

OpenStudy (anonymous):

And I would have to go up to 6.

OpenStudy (anonymous):

\[-2x+2y > -5\] \[2y > 2x + 5\] \[y > x + \frac{ 5 }{ 2 }\]

jimthompson5910 (jim_thompson5910):

to graph this, you can use the points (0, 2.5) and (1, 3.5) draw a line through those two points then make the line a dashed line finally, shade above this dashed line to complete graphing the inequality

OpenStudy (anonymous):

But don't you go up 5, over 2? Instead of 2.5

jimthompson5910 (jim_thompson5910):

5/2 = 2.5 so that's how I got that first point

jimthompson5910 (jim_thompson5910):

you plug in x = 0 to get y = 2.5

OpenStudy (anonymous):

I think I might've just done my whole paper wrong.

jimthompson5910 (jim_thompson5910):

if it was 5/2x then yes you would go up 5 and over 2

jimthompson5910 (jim_thompson5910):

but it's really 1x or 1/1x so you go up 1 and over 1

OpenStudy (anonymous):

Oh ok, thanks for clarifying. Will definetly help me on tomorrows exam. Gotta redo the paper though. :P

jimthompson5910 (jim_thompson5910):

np